### v1

# Beyond Diagonal RIS for 6G Non-Terrestrial Networks: Potentials and Challenges

Abstract
Reconfigurable intelligent surface (RIS) has emerged as a promising technology in both terrestrial and non-terrestrial networks (NTNs) due to its ability to manipulate wireless environments for better connectivity. Significant studies have been focused on conventional RIS with diagonal phase response matrices. This simple RIS architecture, though less expensive, has limited flexibility in engineering the wireless channels. As the latest member of RIS technology, beyond diagonal RIS (BD-RIS) has recently been proposed in terrestrial setups. Due to the interconnected phase response elements, BD-RIS significantly enhances the control over the wireless environment. This work proposes the potential and challenges of BD-RIS in NTNs. We begin with the motivation and recent advances in BD-RIS. Subsequently, we discuss the fundamentals of BD-RIS and NTNs. We then outline the application of BD-RIS in NTNs, followed by a case study on BD-RIS enabled non-orthogonal multiple access low earth orbit satellite communication. Finally, we highlight challenges and research directions with concluding remarks.
I Introduction
Fifth-generation wireless networks have primarily focused on optimizing transceiver designs to manage the uncontrollable wireless environment. However, as we move into the realm of sixth-generation (6G) and beyond networks, there is a shift towards new emerging concepts such as non-terrestrial networks (NTNs) and the coexistence of terrestrial and non-terrestrial networks. While this integration offers numerous benefits, it also presents the challenge of interference when these networks share the same spectrum resources [1]. To address this issue, one possible concept might be manipulating the wireless environment itself, thanks to the innovative technology of reconfigurable intelligent surfaces (RIS). RIS consists of reconfigurable elements capable of altering the propagation environment to enhance the channel gain of the received signals, thus enhancing the spectral efficiency of wireless communications. While the advantages of conventional diagonal RIS have been validated in various applications, it relies on a simple diagonal phase response matrix, where each element independently controls the phase of the incident signal without any cooperation with other elements [2]. This approach has some shortcomings: it only allows phase control, limiting passive beamforming capabilities, and it confines signal reflection to a single direction, thus limiting wireless coverage.
To address these limitations, a new member of RIS technology called beyond diagonal RIS (BD-RIS) has recently merged, which introduces connections among phase response elements that significantly enhance control over the reflected signals [3]. Moreover, the mathematical model for BD-RIS moves beyond the constraints of diagonal matrices, allowing for more sophisticated wave manipulations. Depending on how the phase response elements are connected, various BD-RIS architectures can be classified as single-connected, fully-connected, and group-connected BD-RIS, respectively. This leads to different characteristics of phase response matrices, i.e., block diagonal matrix for group-connected, unitary matrix for fully-connected, and diagonal for single-connected architecture. Note that single-connected is a special case of group-connected such that each group consists of a single phase response element. Fig. 1 envision a broad range of BD-RIS uses in 6G NTNs.
Figure 1: BD-RIS enabled 6G NTNs.
I-A Related Works
Recently, some works have investigated the performance of BD-RIS in 6G wireless networks. In [4], Li et al. have proposed channel estimation and beamforming design for multi-antenna BD-RIS enabled wireless systems to investigate mean square error, spectral efficiency and sum rate. In [5], the authors have derived a closed-form solution for received channel gain maximization of BD-RIS enabled multi-user system by optimizing the active and passive beamforming. In [6], Esmaeilbeig et al. have adopted alternating optimization to maximize the signal-to-noise ratio in BD-RIS enabled integrated sensing and communication systems. In [7], Li et al. have computed the scaling law for received signal and maximized the sum rate of multi-user networks by designing an efficient algorithm for multi-sector BD-RIS. In [8], Mahmood et al. have optimized communication and computational resources to reduce the system latency in BD-RIS equipped unmanned aerial vehicle based mobile edge computing. In [9], Nerini et al. have derived a closed-form solution for designing efficient precoding at the transmitter, combining at the receiver, and beamforming at BD-RIS to maximize the average received sum power of the system. In [10], Soleymani et al. have optimized transmit beamforming and phase response to improve the performance of BD-RIS enabled multi-cell rate splitting multiple access networks. In [11], Liu et al. have jointly designed linear filter, transmit beamforming, and phase response to maximize the sum rate of BD-RIS enabled integrated sensing and communication. Moreover, Zhou et al. [12] have minimized the power consumption and maximized energy efficiency in BD-RIS enabled wireless networks by optimizing the transmit beamforming and BD-RIS phase response. In [13], the authors have jointly optimized the transmit precoder and phase response to maximize the sum rate of BD-RIS enabled multi-user systems. Table I provides the detailed summary of the above works on BD-RIS enabled 6G wireless networks.
Ref. Year Scenario Objective Findings Proposed solution
[4] 2024 Terrestrial CE and BF design Mean square error, SE, SR Used least square method for CE and efficient algorithm for beamforming design.
[5] 2024 Terrestrial Active and passive BF design SR, effective channel gain and time complexity Exploited two-stage low-complex joint active and passive beamforming at BS and BD-RIS.
[6] 2024 Terrestrial Phase response design for BD-RIS Signal-to-noise ratio Adopted an alternating optimization to jointly optimize the $\Phi$ for integrated sensing and communication.
[7] 2023 Terrestrial BF design for multi-sector BD-RIS SR Derived the scaling law for received signal and Designed an efficient algorithm for multi-sector BD-RIS.
[8] 2024 Terrestrial BD-RIS deployment, computational and communication resources Latency, max-min rate, task segmentation and resource allocation Proposed a joint iterative optimization approach for BD-RIS deployment, computational and communication resources.
[9] 2024 Terrestrial Precoding, combining, and BF design Average received sum power and computational complexity Derived a closed-form solution for precoding at transmitter, combining at receiver and BF at BD-RIS.
[10] 2024 Terrestrial BF and phase response design Fairness rate and fairness EE Proposed majorization minimization and alternating optimization to iteratively solve transmit BF and phase response problem for fairness rate and EE improvement.
[11] 2024 Terrestrial Linear filter, transmit BF and BD-RIS phase response design SR Used majorization-minimization and penalty-dual decomposition methods to solve the problem of joint linear filter, transmit BF, and phase response design.
[12] 2023 Terrestrial Transmit BF and BD-RIS phase response design SR, EE, power consumption Utilized a uni-fold optimization approach to solve the problem of total power consumption minimization and EE maximization.
[13] 2023 Terrestrial Transmit precoding and BD-RIS phase response design SR Employed fractional programming theory to solve the transmit precoding and phase response problem for sum rate maximization.
Ref-Reference, CE-Channel estimation, BF-Beamforming, SE-Spectral efficiency, EE-Energy efficiency, SR-Sum rate.
TABLE I: Summary of different works on BD-IRS enabled 6G wireless networks.
Although some research has been done on BD-RIS enabled terrestrial networks, there is no existing literature that considers BD-RIS in NTNs. Motivated by this, we study the potential and challenges of BD-RIS in 6G NTNs. The article first discusses the fundamentals of NTNs and BD-RIS technology. Then, it explains the applications of BD-RIS in NTNs. Subsequently, it provides a case study on BD-RIS enabled low earth orbit (LEO) satellite communication using downlink non-orthogonal multiple access (NOMA). Finally, this work highlights challenges and future research directions in BD-RIS enabled 6G NTNs.
II Non-Terrestrial Networks
NTNs is a sophisticated and diverse network that consists of satellite platforms and aerial platforms. It combines the wide coverage of satellite networks in the space layer and the adaptability of aerial networks in the atmospheric layer, as illustrated in Fig. 1.
II-A Atmospheric Layer
The aerial network in the atmospheric layer primarily consists of UAVs and HAPSs [1]. The UAVs function at varying heights, spanning from hundreds to thousands of meters, and have the capability to cover distances of many kilometers. The operation is straightforward, the deployment is convenient, and the mobility is versatile. It has the capability to establish network connections and transmit signals through relays. However, the device’s operational duration varies from a few minutes to a few hours due to its limited battery capacity. On the other hand, HAPSs are situated in the stratosphere at an altitude ranging from 20 km to 50 km above the Earth’s surface. These platforms tend to remain in a fixed position. A network is established among many HAPSs by mmWave or THz connections. Because HAPSs are positioned at high altitudes, each HAPS has an extensive coverage area that can span up to 100 kilometers in circumference when the elevation angle is 10 degrees. Consequently, by reducing the number of HAPSs, it is possible to obtain a broader coverage area and expedite deployment. HAPS communication systems possess the attributes of cost-effectiveness, minimal latency, rapid deployment, and high capacity in contrast to satellite systems.
The aerial network possesses exceptional flexibility and scalability, making it suitable for a wide range of applications such as emergency communication, network traffic offloading, and temporary communication. For instance, in the event of natural disasters causing the destruction of the terrestrial network, the aerial network can be swiftly deployed to restore connection. Furthermore, an aerial network exhibits a reduced expenditure and decreased delay compared to a satellite network. Nevertheless, the limitations on energy consumption and weight have resulted in a significant drawback for the aerial network, namely its limited endurance period. Moreover, due to the HAPSs restricted height, it can easily cause significant interference to the terrestrial network.
II-B Space Layer
This layer includes different satellite networks. Based on height, the satellite network has three tiers: GEO, MEO, and LEO [1]. GEO satellites form the core space-based network, whereas MEO and LEO satellites form the access network that connects to the aerial or terrestrial network. GEO satellites are set at 35,786 kilometers and the transmission require 0.27 seconds to reach from GEO to ground terminal. GEO satellites cover up to one-third of Earth’s surface. Except for the South and North Poles, three GEO satellites evenly spaced offer global communication. The majority of MEO satellites operate at 8000 miles and they require 0.1 second for signal transmission to the ground terminal. These satellites mostly deliver high-speed data and messages. Many satellites operate exclusively in this orbit for telecommunications and military use. LEO satellites orbit 500–1000 miles above sea level. From this orbit, the satellite’s signal reaches the ground station in 0.05 seconds. This track is best for low latency applications.
Thanks to their exceptionally high altitude, a limited number of satellites can provide remarkably wide coverage, unaffected by geographical features. This is particularly crucial for rural regions and maritime communication. Simultaneously, satellites harness solar panels to generate energy, with an average lifespan of approximately 10 years, making them a superb addition to terrestrial communication. These advantages facilitate the growth of satellite networks. As an illustration, in 2023, the United States successfully deployed Viasat-3 satellites with a remarkable capacity of 1 terabits per second (Tbps). Additionally, SpaceX has launched a multitude of LEO Starlink satellites. Nevertheless, the existing satellite network continues to suffer from limitations in transmission capacity and significant latency.
III Modes and Advantages of BD-RIS in NTNs
In this section, we first study the different modes BD-RIS technology and then we discuss its advantages compared to the conventional diagonal RIS technology.
III-A Modes of BD-RIS in NTNs
BD-RIS can be classified into three modes, including reflective BD-RIS, hybrid BD-RIS, and multi-sector BD-RIS [13].
III-A1 Reflective BD-RIS
In this mode, the BD-RIS reflects signals that come from one side back towards that side, resulting in coverage of 180-degree space. As can be seen in the satellite communication scenario in Fig. 2(i), all of the reflecting elements of the BD-RIS are oriented in the same direction in order to enable the reflective mode during transmission. When considering the BD-RIS with reflecting mode, the phase shift matrix $\boldsymbol{\Phi}$, which is subject to a unitary constraint, stands out as the defining characteristic. The matrix $\boldsymbol{\Phi}$ is unitary if it satisfies the following condition: $\boldsymbol{\Phi}^{*}\boldsymbol{\Phi}=\boldsymbol{\Phi}\boldsymbol{\Phi}^{*}={\bf I}$, where $\boldsymbol{\Phi}^{*}$ is the conjugate transpose of $\boldsymbol{\Phi}$ and ${\bf I}$ denotes an identity matrix. This property ensures that the columns (and rows) of the matrix are orthogonal. In the context of BD-RIS for 6G wireless networks, unitary constraints are applied to the design of these surfaces to ensure efficient signal manipulation without loss of power. For example, the BD-RIS may be designed to apply unitary transformations to incoming signals, maintaining signal integrity and improving communication performance by optimizing reflection and transmission paths.
III-A2 Hybrid BD-RIS
This mode allows signals that are impinging on one side of the BD-RIS to be partially reflected toward the same side and partially transmitted toward the opposing side, which results in a coverage of 360-degree space. The BD-RIS that operates in hybrid mode is also referred to as the intelligent omni-surface or the simultaneous transmitting and reflecting BD-RIS. In order to support the hybrid mode, two elements with a unidirectional radiation pattern are placed back to back. These elements are then connected to a two-port fully connected reconfigurable impedance network, as shown in HAPS communication scenario in Fig. 2(ii). This ensures that each reflecting element and each transmitting element cover 180-degree space, respectively, in order to achieve 360-degree coverage. In terms of mathematics, the BD-RIS with $K$ number of phase response elements and operating in hybrid mode is characterized by two matrices, namely transmitting and reflecting, i.e., $\boldsymbol{\Phi}_{t}\in\mathbb{C}^{K/2\times K/2}$ and $\boldsymbol{\Phi}_{r}\in\mathbb{C}^{K/2\times K/2}$, where $\boldsymbol{\Phi}_{t}$ is a matrix of transmitting phase response elements and $\boldsymbol{\Phi}_{r}$ denotes the matrix of reflecting phase response elements. These matrices should satisfy $\boldsymbol{\Phi}^{H}_{t}\boldsymbol{\Phi}_{t}+\boldsymbol{\Phi}^{H}_{r}\boldsymbol{\Phi}_{r}={\bf I}_{K/2}$.
III-A3 Multi-Sector BD-RIS
This mode is an extension or a more comprehensive version of the hybrid mode. In multi-sector BD-RIS, the entire area is partitioned into $S$ sectors, where $S$ must be greater than or equal to 2. Signals that come into contact with one sector of the BD-RIS can be partially redirected back to the same sector and partially dispersed toward the remaining $S-1$ sectors. In order to facilitate the multi-sector mode of BD-RIS, each sector consists of $S$ phase response elements positioned at the edges of an $S$-sided polygon. These elements have a unidirectional radiation pattern that covers the $1/2$ area of the sector, preventing any overlap between adjacent sectors. The $S$ elements are connected to a $S$-port fully connected reconfigurable impedance network, as illustrated in UAV communication scenario in 2(iii). Therefore, the multi-sector mode of BD-RIS can cover 360-degree space, similar to the hybrid BD-RIS. However, multi-sector BD-RIS offers greater performance improvements compared to the hybrid mode due to the utilization of high-gain elements with narrower beamwidth, which covers 180-degree space. The BD-RIS with multi-sector mode can be mathematically described by a set of $S$ matrices, $\boldsymbol{\Phi}_{s}\in\mathbb{C}^{K/S\times K/S}$, $s\in S$. These matrices should meet the condition of $\sum_{s=1}^{S}\boldsymbol{\Phi}_{s}^{H}\boldsymbol{\Phi}_{s}={\bf I}_{K/S}$.
Figure 2: Different modes of BD-RIS.
III-B Advantages of BD-RIS in NTNs
BD-RIS provides substantial benefits compared to the traditional diagonal RIS system due to the connection between elements and the full scattering matrix. BD-RIS offers advanced beamforming capabilities, superior signal manipulation, enhanced flexibility and adaptability, increased spectrum efficiency, improved interference management, and greater resilience to environmental changes. These advances greatly enhance the capabilities of BD-IRS, making it a highly potent and adaptable technology for improving future wireless communication networks, particularly those designed for 6G and beyond. Key advantages of BD-RIS over diagonal RIS can be further discussed in the following.
III-B1 Advance Beamforming
The functionality of conventional RIS is restricted to controlling the phase of the incident signal. As a result, its capacity to manipulate passive beamforming is limited, leading to a degradation in performance. BD-RIS provides inter-element connections, however, this comes at the cost of increased circuit complexity; its mathematical description is not restricted to diagonal phase response matrices. This makes for more precise and adaptive beamforming, which boosts signal quality and strength in many directions.
III-B2 Improved Signal Manipulation
In contrast to the conventional diagonal RIS, which can only manipulate the diagonal elements of the phase response matrix, BD-RIS comparatively offers greater flexibility by allowing manipulation of both diagonal and non-diagonal elements of the phase response matrix. This enhanced capability significantly improves the performance of various wireless communication scenarios in 6G networks.
III-B3 Enhanced Flexibility and Adaptability
Conventional RIS is efficient in situations where signal paths are more consistent and can be predicted. BD-RIS, in contrast, offers enhanced adaptability to accommodate fluctuating signal conditions, including movements in user positions, mobile blockages, and evolving network requirements. This flexibility guarantees constant performance even under extremely unpredictable settings.
III-B4 Better Interference Management
Diagonal RIS can reduce interference by adjusting the phase of incident signals. BD-IRS offers exceptional interference management capabilities through the utilization of modern signal processing techniques. This leads to reduced levels of interference, hence improving the overall performance of the network and the experience of the user.
III-B5 360-Degree Coverage
Unlike diagonal RIS, which has a limited coverage of 180-degrees, the BD-RIS operates in hybrid and multi-sector modes, providing 360-degree coverage. Furthermore, the multi-sector mode can enhance channel gain by utilizing smaller beamwidth and higher element gain of each BD-RIS element. This effectively extends the communication range, ensuring full-space coverage.
IV Applications of BD-RIS in NTNs
Well-established research on conventional diagonal RIS, clearly demonstrates its use cases and applications in diverse fields, including, but not limited to, wireless power transfer, coverage improvements, and wireless sensing. In the following sections, we showcase, for the first time, some of the potential applications of BD-RIS in an NTN-enabled communication scenario.
IV-A Seamless and Ultra Precision Indoor Positioning
It is well-known that wireless signals based on NTN are unable to penetrate indoor environments effectively, as they are weakened and scattered by roofs, walls, and other structures. Moreover, the availability of line-of-sight (LoS) signals greatly impacts positioning accuracy. These limitations necessitate a separate infrastructure for indoor positioning applications, thereby increasing overall costs and energy requirements. Since BD-RIS can flexibly reflect signals in multiple controlled directions, it can play a pivotal role in providing an effective conduit for NTN signals to overcome penetration losses caused by building structures. Furthermore, compared to conventional diagonal RIS, BD-RIS offers enhanced beamforming capabilities, thus improving a network’s ability to achieve ultra-precise localization of devices. In conclusion, these additional features of BD-RIS can be crucial in developing a new class of indoor positioning systems that do not require any additional infrastructure, costs, or energy resources.
IV-B Improved IoT Connectivity
The Internet of Things (IoT) has been a familiar concept for quite some time, but its popularity surged with the integration of machine-type communication (mMTC) as a feature of 5G. IoT encompasses a diverse range of applications, spanning from smart home automation to monitoring deforestation activities. Particularly in areas with limited or no cellular coverage, connecting these IoT devices to the internet poses a significant challenge. To address such obstacles, the 3GPP release 17 incorporates NTN into traditional Terrestrial Networks (TN), enabling NTNs to serve as a bridge between the internet and these IoT devices. Despite playing a critical role in overcoming these limitations, NTN introduces its own set of challenges, such as coverage (while considering smart home indoor environment applications in remote/rural areas) and power consumption. Similarly, the enhanced capabilities of BD-RIS, which include reflecting signals in multiple controlled directions and improved beamforming, have the potential to mitigate these barriers and establish a seamless connection between NTN and IoT devices.
IV-C Extended Coverage for NTN-aided Terrestrial Users
One of the core applications of conventional diagonal RIS, for which it was originally developed, is to provide extended coverage, overcome blind spots, and ensure seamless connectivity for terrestrial users. In the context of NTN-aided communication, conventional diagonal RIS has shown great potential in improving the performance of cellular users with weak or no connectivity to the base station. For instance, as demonstrated in [14], conventional diagonal RIS-aided UAVs (a specific use case of NTN) have clearly illustrated the usefulness of conventional diagonal RIS in NTN systems. Although conventional diagonal RIS-aided NTN has shown significant potential in enhancing the performance of terrestrial users, the limitations associated with conventional diagonal RIS prevent us from fully benefiting from NTN. As discussed in the preceding sections, BD-RIS clearly surpasses conventional diagonal RIS by offering greater flexibility in reflecting signals in multiple controlled directions. This capability provides more extended coverage compared to conventional diagonal RIS and plays a significant role in increasing the performance of NTN.
Figure 3: BD-RIS enabled LEO communication.
V Case Study: BD-RIS Enabled NOMA LEO Satellite Communications
In this section, we provide a case study on BD-RIS enabled NTNs. We will first discuss the considered system model and proposed optimization. Then, we provide the numerical results and their discussion.
V-A System Model and Proposed Optimization
Let us consider a downlink scenario of BD-RIS enabled satellite communication, as shown in Fig. 3, where an LEO communicates with ground users (GUs) using the power domain NOMA. In practice, the GUs operate with low directional antennas, so closing links between LEO and GUs is challenging when the network operates on a high frequency. Moreover, the direct links between LEO and GUs might not be possible, considering the mobile scenario. To compensate for path losses due to poor link budget and improve the received signal strength, we consider BD-RIS installed on the optimal location and assisting the signal delivery from LEO to GUs. Thus, GUs can receive signals from direct and BD-RIS enabled links. Assuming perfect acquisition of channel state information (CSI) by channel estimation techniques, this scenario provides an upper-bound solution.
This case study seeks to improve the achievable sum-rate of BD-RIS enabled NOMA LEO satellite communication by optimizing the transmission power of LEO and phase response of BD-RIS while ensuring the minimum data rate requirements of GUs. This is achieved by formulating a mathematical optimization problem for achievable sum-rate maximization which is subjected to several practical constraints such as minimum rate constraint, LEO power budget constraint, NOMA power allocation constraint, and BD-RIS phase response constraint, respectively. The formulated optimization is nonlinear and NP-hard in nature due to coupling decision variables and NOMA interference. In practice, solving such problems is challenging due to high computational complexity. To make the optimization tractable, we first adopt the block descent coordinate method to split the joint optimization into two problems: 1) transmission power at LEO for a given phase response at BD-RIS, and 2) phase response at BD-RIS, given the power allocation for LEO. Then, each problem is further transformed into linear form and simultaneously solved using a standard convex optimization such as CVX toolbox.
Figure 4: Performance Analysis
V-B Numerical Results and Discussion
This section presents the numerical results to reveal the effectiveness of the proposed LEO communication scheme in the efficient allocation of NOMA transmission power and phase response of the BD-RIS. The analysis is carried out by setting the transmit power to 20 dBm, the operating frequency to 3.5 GHz, and the elevation angles to $45^{\circ}$. The antenna gains are set to 10 dBi, with a path-loss exponent of 2.5. The dimension of the BD-RIS elements is $0.5\,\text{m}\times 0.5\,\text{m}$, and the distance from BD-RIS to the satellite is 500 km. The altitude of the satellite is set to 600 km, and the radius of Earth is considered to be 6371 km. The reflection coefficient magnitude is 0.9, the noise power is -90 dBm, and the Rician factor is set as 10.
The simulations were conducted based on Monte Carlo simulations, where the average results were obtained over independent channel realizations. Additionally, the simulation results were plotted considering the sum rate across different numbers of BD-RIS elements. Furthermore, the results were compared with relevant benchmark schemes, including the optimal BD-IRS scheme with orthogonal multiple access (OMA) technique and NOMA scheme with random BD-RIS phase response. Result in Fig.4 shows that as the number of BD-RIS elements increases, the sum rate of the system also increases. This is because the BD-RIS effectively directs the signal toward the specific user, resulting in an increase in channel gain. However, the proposed optimal BD-RIS scheme achieves high sum rate compared to the relevent benchmark schemes. The main reason of this is joint power allocation and phase response design of the proposed BD-RIS enabled NOMA LEO communication scheme.
VI Challenges and Research Directions
Although the information provided in the previous sections well establishes that BD-RIS offers significant benefits over conventional diagonal RIS, it comes with limitations that open up crucial problems needing resolution to fully exploit its advantages. On one hand, there is very limited literature highlighting the key issues associated with the use of BD-RIS in actual wireless networks and providing new research directions. On the other hand, since the use of NTN for terrestrial users is still in its infancy, the issues related to the use of BD-RIS in such settings require extensive analysis. Therefore, in this paper, we have highlighted three key issues associated with the use of BD-RIS in an NTN-assisted wireless network.
VI-A Adaptive Channel Realization
The accurate realization of channels is one of the key aspects that significantly supports BD-RIS in achieving its promised performance gains. The conventional channel estimation techniques used in conventional diagonal RIS, such as semi-passive methods (enabling conventional diagonal RIS to perform pilot transmission using low-power RF chains), can easily be replicated in BD-RIS at the expense of higher power consumption. However, purely passive techniques cannot be directly adopted due to variations in RF circuitry. When incorporating BD-RIS in a NTN wireless communication scenario, the channel estimation process becomes much more complicated as the signal propagation path significantly lengthens, and fast time-varying channels are introduced (considering the LEO satellite use case of NTN). Therefore, the channel estimation techniques used in conventional diagonal RIS cannot be directly adapted to the BD-RIS-assisted NTN scenario. Instead, innovative upgrades and/or new techniques are required that not only improve the accuracy of channel estimation but also support low power consumption, promoting sustainability.
VI-B RF Circuitry Constraints
The development of BD-RIS is still in its early stages, where the mathematical models used to simulate it for various applications mostly rely on ideal assumptions, such as lossless impedance components, perfect matching, and no mutual coupling between BD-RIS elements. These assumptions directly impact the beamforming capability of BD-RIS by increasing the complexity of performing beamforming. Moreover, they also limit the implementation of the BD-RIS matrix using discrete values, while relying on simple quantization mechanisms. Therefore, it is necessary to consider these impairments when designing BD-RIS. For BD-RIS-assisted NTN applications, these RF circuitry constraints present additional challenges. Signals from NTN systems travel significant distances and face atmospheric turbulence, so relying on ideal models in two-way communication significantly impacts BD-RIS-assisted NTN communication. To fully exploit the benefits of BD-RIS in NTN-assisted scenarios, RF circuitry constraints need to be fully incorporated into mathematical models and simulation analyses using innovative methods.
VI-C Receiver Sensitivity
One of the main concerns in NTN-assisted terrestrial users is the distance traveled by the wireless signal between the transmitter and the receiver. Although there is supposed to be an acceptable quality of LoS connection between a satellite and a ground user, the free space path loss and other atmospheric turbulence significantly degrade the signal quality. To fully exploit the benefits of BD-RIS in an NTN scenario, the sensitivity of BD-RIS needs to be effective enough to differentiate between noise and the NTN signal. In conventional diagonal RIS, significant attention has been given to overcoming this issue. For instance, in [15], the structure of conventional diagonal RIS is repurposed to address angular sensitivity by utilizing metallic vias to develop programmable meta-atoms that are insensitive to angular distortions. A similar setup can be adopted for BD-RIS, but when considering NTN-assisted BD-RIS wireless communication scenarios, this issue needs to be addressed with innovative designs and algorithms.
VII Conclusion
This article proposed BD-RIS for 6G NTNs to further enhance control over the radio environment with a beyond diagonal phase response matrices. First, we provided the fundamentals of 6G NTNs and discussed the atmospheric layer and space layer, followed by various modes and advantages of BD-RIS technology. Then, we discussed different applications of BD-RIS in 6G NTNs. To investigate the performance of BD-RIS in NTNs, we also provided a case study on BD-RIS enabled NOMA LEO satellite communication. Finally, we studied challenges and future research directions in BD-RIS enabled NTNs and concluded our work.